OpenStudy (anonymous):

roots -2 and 4 and vertex 1,-10

OpenStudy (anonymous):

Is this a parabola?

OpenStudy (anonymous):

quadratic function

OpenStudy (anonymous):

Ah, same thing. The general form of a quadratic function is \(y=ax^2+bx+c\). You're given that the roots are \(-2\) and \(4\), which means \(y=0\) when \(x=-2,4\). So you have the following \[0=a(-2)^2+b(-2)+c\\ 0=a(4)^2+b(4)+c\] \[\begin{cases}4a-2b+c=0\\16a+4b+c=0\end{cases}\] You're also given another point, \((1,-10)\), which means \(y=-10\) when \(x=1\). Plugging in this \(x\), you get \[-10=a(1)^2+b(1)+c\\ a+b+c=-10\] So, in sum, you have the following system of three equations with three unknowns: \[\begin{cases}4a-2b+c=0\\16a+4b+c=0\\a+b+c=-10 \end{cases}\] Your task now is to solve for \(a,b,c\).